You receive a loan for $1000. The loan accumulates interest at a rate of 6.5%. How much interest will you owe in 4 yrs?

a) 286.47 b)1286.47 c)260 d)1260 e)1000
Answer #1

FV = PV(1+R)^N

FV= 1000(1+0.065)^4

FV= 1000(1.065)^4

FV=1000*1.286466

FV = 1286.466

INTEREST = 1286.466-1000

INTEREST = 286.466

SO OPTION A 286.47 IS CORRECT.